OpenStudy (anonymous):

Establish each identity: 1-2cos^2 Θ/sinΘ cosΘ = tan Θ - cotΘ

OpenStudy (anonymous):

We know: \[ 2\cos^2(\theta)-1=\cos(2\theta) \]So: \[ 1-2\cos^2(\theta)=-(2\cos^2(\theta)-1)=-\cos(2\theta)=\sin^2(\theta)-cos^2(\theta) \]Try using that.